[Brahms-l] Shift report 20040102 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Fri Jan 02 2004 - 15:56:29 EST
------------------------------------------------------------
Date: Fri Jan  2 15:56:28 2004, Shift: 08:00-16:00
Supervisor: JHL, with: 
Log book#15 pages 21 - 22

Shift Summary:
------------------------------------------------------------
1. Beam until ~10 am (a long store ~10hr)
2. Data
   FS@8deg 1/3B MRS@90deg 6B
   run 9316: 290K trigger1
3. Problems of having events with 0 adc in ZDC 
   has been fixed (FV: See the logbook for details)
   No new problems seen

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jan 2 15:57:09 2004

This archive was generated by hypermail 2.1.8 : Fri Jan 02 2004 - 15:57:23 EST